In a ten-question multiple-choice test with four choices for each question, a student who was not prepared guesses on each item. Find each probability.
a. P(six questions correct)
b. P(at least eight questions are correct)
c. P(fewer than eight questions correct)

Answer by stanbon(75887)

In a ten-question multiple-choice test with four choices for each question, a student who was not prepared guesses on each item. Find each probability.
P(correct) = 1/4 ; P(wrong) = 3/4
--------------------------------------------
a. P(six questions correct)= 10C6*(1/4)^6*(3/4)^4 = 0.01622
-------------------------------------------------
b. P(at least eight questions are correct)= 1-binomcdf(10,1/4,7) = 0.000415802...
-------------------------------------------------
c. P(fewer than eight questions correct) binomcdf(10,1/4,7) = 0.99958
========================================================================
Cheers,
Stan H.